i Requirements 1. 2. Compute Creatives days sales in receivables for 2018. (Round to the nearest day.) Suppose Creatives nCreative Media Sign Incorporated sells on account. Recently. Creative reported the following figures: $ Net Credit Sales Net
Answer #1
Statement showing Computations
Particulars Amount
Net receivables at beginning          49,100.00
Net receivables at ending          38,900.00
Average receivables = (49100 + 38900)/2          44,000.00
Net credit sales         580,800.00
Average accounts receivable turnover = 580,800/44000                  13.20
Days sales in receivable = 365/13.20                        28
2)
Company's collection period is good.
2/10 net 30 means if payment is made withing 10 days then 2% discount is given, otherwise payment should be made within 30 days.
In above case days sale in receivable is 28 days which is less than 30 days
Thus collection period is good